# Small Language Model Can Self-correct

### v1

@@ Abstract
Generative Language Models (LMs) such as ChatGPT have exhibited remarkable performance across various downstream tasks. Nevertheless, one of their most prominent drawbacks is generating inaccurate or false information with a confident tone. Previous studies have devised sophisticated pipelines and prompts to induce large LMs to exhibit the capability for self-correction. However, large LMs are explicitly prompted to verify and modify its answers separately rather than completing all steps spontaneously like humans. Moreover, these complex prompts are extremely challenging for small LMs to follow. In this paper, we introduce the Intrinsic Self-Correction (ISC) in generative language models, aiming to correct the initial output of LMs in a self-triggered manner, even for those small LMs with 6 billion parameters. Specifically, we devise a pipeline for constructing self-correction data and propose Partial Answer Masking (PAM), aiming to endow the model with the capability for intrinsic self-correction through fine-tuning. We conduct experiments using LMs with parameters sizes ranging from 6 billion to 13 billion in two tasks, including commonsense reasoning and factual knowledge reasoning. Our experiments demonstrate that the outputs generated using ISC outperform those generated without self-correction. We believe that the output quality of even small LMs can be further improved by empowering them with the ability to intrinsic self-correct.
@@ Introduction
Generative Language Models (LMs) have gained considerable attention due to their remarkable capabilities [4; 5]. Despite the convincing and realistic nature of text generated by these LMs, a concern with LMs lies in their tendency to produce fabricated facts and generate false information [6]. Moreover, these models deliver inaccurate information employing unequivocal expressions, which poses substantial risks as it can lead to the spread of misleading and harmful content.
@@
One of the contributing factors to the hallucination lies in inadequate acquisition of knowledge [8; 9]. For example, consider the question Which animal is China’s nation treasure?, LMs may provide a different animal name like tiger instead of panda due to a lack of relevant knowledge. Considerable efforts have been made to alleviate such hallucination induced by lacking of knowledge in LMs. One approach involves supervised fine-tuning LMs with standard ground-truth answers to enhance their comprehension of relevant knowledge [10; 11]. This method has shown promising efficacy. However, it demands a significant amount of high-quality annotated data for training. Additionally, other methods have relied on external verifier or critic model to evaluate the accuracy of a statement [18; 19]. Training a verifier necessitates a large number of high-quality evaluation annotations and further fine-tuning of the model, which restricts its broad applicability to other tasks and domains.
@@
Another reason for an LM to provide incorrect response is intrinsically linked to the design architecture of generative language models themselves [14; 19; 21]. It is widely acknowledged that LMs generate a sentence by maximizing the likelihood of the next token given all previous tokens. Subtle differences in the preceding sentences can potentially lead to diverse generation outcomes. For example, when the question is Who is the author of The Analects?, the model gives the correct answer as Confucius. However, when the input question becomes Is Laozi or Confucius the author of the Analects of Confucius?, the model is likely to generate an answer of Laozi. In this case, the model has the ability to rely on its knowledge to recognize false information [16]. This process is akin to how humans perform self-verification of answers to minimize mistakes [17]. Moreover, when we realize our answer is wrong, we further modify it. Motivated by this, when the model itself detects the potential hallucination, the next step is to correct the error or mistake. Once the model incorporates this inherent self-correction mechanism, it can address similar issues in other domains and achieve self-improvement.
@@
The existing work [20; 29] towards self-correction in LMs has mainly focused on lager models like ChatGPT and GPT4, which is challenging to migrate these self-correction methods to small LMs. Some studies indicated that the self-correction ability depends on model parameters and only emerges in models with larger parameter sizes [14]. The main reason is that they devised a sophisticated pipeline and zero-shot prompts to achieve self-correction. However, these prompts crafted for self-verification and self-modification are difficult for small models to understand. As depicted in Figure 1, upon generating the initial answer to the given question, an additional feedback instruction is utilized to guide ChatGPT in generating feedback information regarding the initial answer. This information contains an evaluation of the correctness of the initial answer. A subsequent modification instruction is employed to alter or refine the initial answer based on the feedback received.
@@
Nevertheless, small models typically lack self-awareness [3] and tend to exhibit greater confidence in their generated responses. Consequently, they struggle to assess the quality of their generated outcomes. The capability for self-verification serves as a prerequisite for achieving self-correction. Furthermore, the manner in which self-correction is achieved through multi-step prompt engineering within LMs differs from the spontaneous and one-time correction observed in humans.
@@
  • To the best of our knowledge, we are the first to demonstrate that small language models with even 6 billion parameters possess the capacity for self-correction during response generation without relying ground truth.
@@
Instruction fine-tuning LMs often struggle to maintain alignment well with human intent after pre-training. However, this issue can be alleviated by employing techniques such as Instruction Fine-Tuning (IFT) [23; 11] and Reinforcement Learning from Human or AI Feedback (RLHF/RLAIF) [24; 30]. IFT involves training an LM to generate responses based on provided prompts, which might optionally include task instructions. The data used for fine-tuning is usually less than the original pre-training corpus. These datasets used in IFT are often curated through crowd-sourcing or extracted from an LM that is already capable of generating instruction-following examples. Studies have shown that IFT can enhance zero-shot generalization to various unseen tasks. To optimize the fine-tuning process, various strategies have been proposed to make it more parameter-efficient. These strategies include using adapters [25], prompt tuning [26], and other techniques.
Chain-of thought in language model. Chain-of-Thought (COT), as an alternative prompt strategy, adopts a chained reasoning approach, incorporating a multi-step reasoning path before generating the final answer. This strategy aims to enhance the model’s capacity to handle complex and multi-step queries. kojima2022large [27] introduced a novel yet simple approach to prompt the LM by using the phase Let’s think step by step, which enables reasoning generation in a zero-shot manner. zhou2023leasttomost [13] further decomposed the questions into multiple sub-questions, encouraging the model to engage in a sequential thought process while arriving at the answer. Overall, COT aims to enhance the answer quality through a multi-step reasoning process.
@@
Self-improvement in language model. Richer and more detailed feedback plays a crucial role in aligning the output of the LM with the user’s preferences and in significantly improving the overall performance of the model. Welleck2022GeneratingSB [1] introduced a corrector model used to improve the initial response by learning the specific mistakes made by the LM and the appropriate method for modifying them. Similarly, paul2023refiner [19] introduced a critic model to provide feedback on the LM’s reasoning errors, evaluating the generated intermediate steps. The feedback, together with the original question and preceding intermediate reasoning steps, was subsequently fed back to the LM to improve the generation of the next step.
Scaling model size can increase model’s various capabilities[31], including self-verification (or self-feedback). Recent research efforts have focused on exploring large LMs that employ self-feedback rather than relying on additional models. madaan2023selfrefine [20] introduced an iterative self-refinement algorithm that alternates between feedback and refinement. They guided the model to provide feedback about the initial response by employing few-shot prompts. This feedback is then passed back to the same model to help refine the initial response. Some methods used self-improvement to improve the reasoning ability of LM without supervised data. For example, the LM first generates multiple COT reasoning paths and corresponding answers for each question [2]. The answer with the highest consistency is selected by majority voting. The COT reasoning paths exhibiting high-confidence, along with the selected answers, are augmented by mixed formats to serve as the training data used for IFT. Additionally, the LM verified its multiple generated answers in turn and calculated the verification score based on the number of predicted masked values [3].
@@
($A_{1}^{1}$) Therefore, the correct answer is E. amusement park
@@ Methods
Additionally, we use two methods to enhance the model’s understanding of various self-correction instructions. First, we use gpt-3.5-turbo to generate diverse prompts for different types of instructions mentioned above. We present a brief prompt template below I want you act as a Prompt Rewriter. Your objective is to rewrite a given prompt to make language model to understand. The rewritten prompt must ensure that the requirement remains unchanged. #Given Prompt#:$\left[TaskP\mid COTP\mid COTP\mid SCP\right]$. Second, we randomly select one instruction from TaskP, COTP and SCP sets and combine them. The order of these instructions can also be rearranged, but TaskP is positioned at the beginning.
@@
Answer preparation. To ensure answer diversity, we utilize nucleus sampling [33] to generate multiple answers. After generating multiple answers for a question, the next step is to evaluate the accuracy of each answer by comparing it with the provided ground truth for the each question. In the case of multiple-choice questions, we extract the options of the final answer through string matching and then directly compare them with the standard answer to check the accuracy.
We categorize the generated answers of $M$ into two groups: good cases and bad cases, representing whether $M$ can provide correct answers to the questions, respectively.
@@
In our paper, we apply this loss calculation method to good cases. However, it is not suitable for bad cases due to two reasons. First, these bad cases contain error information, which increases the likelihood of the common issue of hallucination inherent in language models. Second, deliberately training the model to first generate incorrect answers and then correcting them does not align with our intention of teaching the LM how to self-correct. Our aim is instead to enable the model to spontaneously correct itself when it generates information inconsistent with its internal knowledge. Therefore, for bad cases, we refrain from computing the loss corresponding to the incorrect answer part in the output during training. Instead, we calculate the loss from the output on self-verification. It means that for bad cases, only the self-verification and modified correct answer contribute to the loss calculation and parameter update. We call this training method Partial Answer Masking (PAM). In Table 1, the yellow-highlighted part in the output are excluded from the loss calculation.
@@
Base language models. Our goal is to evaluate whether we can improve the performance of any LMs through our proposed ISC, even to small LMs, using ISC. We opt for open-source models with parameters ranging between 6 billion and 13 billion. We use CuteGPT-7B, CuteGPT-13B [^3] https://github.com/Abbey4799/CuteGPT/, ChatGLM-6B [^4] https://github.com/THUDM/ChatGLM-6B, Llama2-7B [^5] https://huggingface.co/meta-llama/Llama-2-7b-hf, Vicuna-7B [^6] https://huggingface.co/lmsys/vicuna-7b-v1.3 and Vicuna-13B [^7] https://huggingface.co/lmsys/vicuna-13b-v1.3 as our instruction fine-tuning base models for continuing training. These models have been fine-tuned on the instruction data and have the ability to follow instructions. CuteGPT is based on the original Llama model structure, expands the Chinese vocabulary and performs pre-training. CuteGPT has two public versions: CuteGPT-7B and CuteGPT-13B. ChatGLM-6B also is an open bilingual language model. It is trained for about 1T tokens of Chinese and English corpus, and demonstrates outstanding performance among language models of equal parameter size. Llama2 is a family of state-of-the-art open-access large language models released by Meta, ranging in scale from 7 billion to 70 billion parameters. Vicuna is also an open-source chatbot trained by fine-tuning Llama.
@@
In our experiments, To assess the capability for self-modification of LM, we use R2R, R2W, W2R, and W2W to denote the times that the LM modifies the correct answer to the right answer, changes the right answer to the wrong answer, changes the wrong answer to another incorrect answer, and modifies the wrong answer to the right answer, respectively.
@@
Discovery 4: By examining the values within the R2R column, we discern that even though our experiments primarily focused on self-verification of the final answers rather than the problem-solving process, we note that the models, through self-correction, could identify inadequacies in the analytical process and subsequently provide supplementary anslysis.
@@
We discover that ISC remains effective for the new task. After the second round of correction, the accuracy of all LMs improve. It is worth noting that for CuteGPT-7B, the ability of self-correction exhibits excellent generalization performance in the new task, with the most significant performance gains.
@@ Conclusion
We introduce Intrinsic Self-Correction (ISC) in LMs, an approach that utilizes models’ own capabilities to identify and further modify their initial responses autonomously. This strong capability can even be applied to smaller LMs. We first devise a general process for constructing self-correction data, applicable to generating diverse self-correction task training data. Furthermore, we introduce a novel fine-tuning method named Partial Answer Masking to instruct LMs to self-correct. We conduct experiments on several open-source LMs to validate the efficacy of ISC. The experimental results on two distinct tasks consistently demonstrate that the utilization of ISC empowers the models with the capability for self-correction, and improves the accuracy of generated answers. In the best case, the accuracy enhancement reaches up to 5.6%.

### v2

@@ Abstract
Generative Language Models (LMs) such as ChatGPT have exhibited remarkable performance across various downstream tasks. Nevertheless, one of their most prominent drawbacks is generating inaccurate or false information with a confident tone. Previous studies have devised sophisticated pipelines and prompts to induce large LMs to exhibit the capability for self-correction. However, large LMs are explicitly prompted to verify and modify their answers separately rather than completing all steps spontaneously like humans. Moreover, these complex prompts are extremely challenging for small LMs to follow. In this paper, we introduce the Intrinsic Self-Correction (ISC) in generative language models, aiming to correct the initial output of LMs in a self-triggered manner, even for those small LMs with 6 billion parameters. Specifically, we devise a pipeline for constructing self-correction data and propose Partial Answer Masking (PAM), aiming to endow the model with the capability for intrinsic self-correction through fine-tuning. We conduct experiments using LMs with parameters sizes ranging from 6 billion to 13 billion in two tasks, including commonsense reasoning and factual knowledge reasoning. Our experiments demonstrate that the outputs generated using ISC outperform those generated without self-correction. We believe that the output quality of even small LMs can be further improved by empowering them with the ability to intrinsic self-correct.
@@ Introduction
Generative Language Models (LMs) have gained considerable attention due to their remarkable capabilities (Guo et al. 2023; Suzgun et al. 2023). Despite the convincing and realistic nature of text generated by these LMs, a concern with LMs lies in their tendency to produce fabricated facts and generate false information (Lin, Hilton, and Evans 2022). Moreover, these models deliver inaccurate information employing unequivocal expressions, which poses substantial risks as it can lead to the spread of misleading and harmful content.
@@
One of the contributing factors to the hallucination lies in inadequate acquisition of knowledge (Manakul, Liusie, and Gales 2023; Huang et al. 2023). For example, consider the question Which animal is China’s nation treasure?, LMs may provide a different animal name like tiger instead of panda due to a lack of relevant knowledge. Considerable efforts have been made to alleviate such hallucination induced by lacking of knowledge in LMs. One approach involves supervised fine-tuning LMs with standard ground-truth answers to enhance their comprehension of relevant knowledge (Wei et al. 2022a; Ouyang et al. 2022). This method has shown promising efficacy. However, it demands a significant amount of high-quality annotated data for training. Additionally, other methods have relied on external verifier or critic model to evaluate the accuracy of a statement (Yang et al. 2022; Paul et al. 2023). Training a verifier necessitates a large number of high-quality evaluation annotations and further fine-tuning of the model, which restricts its broad applicability to other tasks and domains.
@@
Another reason for an LM to provide incorrect response is intrinsically linked to the design architecture of generative language models themselves (Azaria and Mitchell 2023; Paul et al. 2023; Shinn et al. 2023). It is widely acknowledged that LMs generate a sentence by maximizing the likelihood of the next token given all previous tokens. Subtle differences in the preceding sentences can potentially lead to diverse generation outcomes. For example, when the question is Who is the author of The Analects?, the model gives the correct answer as Confucius. However, when the input question becomes Is Laozi or Confucius the author of the Analects of Confucius?, the model is likely to generate an answer of Laozi. In this case, the model has the ability to rely on its knowledge to recognize false information (Schick et al. 2023). This process is akin to how humans perform self-verification of answers to minimize mistakes (Flower and Hayes 1981). Moreover, when we realize our answer is wrong, we further modify it. Motivated by this, when the model itself detects the potential hallucination, the next step is to correct the error or mistake. Once the model incorporates this inherent self-correction mechanism, it can address similar issues in other domains and achieve self-improvement.
@@
The existing work (Madaan et al. 2023; Ganguli et al. 2023) towards self-correction in LMs has mainly focused on lager models like ChatGPT and GPT4, which is challenging to migrate these self-correction methods to small LMs. Some studies indicated that the self-correction ability depends on model parameters and only emerges in models with larger parameter sizes (Azaria and Mitchell 2023). The main reason is that they devised a sophisticated pipeline and zero-shot prompts to achieve self-correction. However, these prompts crafted for self-verification and self-modification are difficult for small models to understand. As shown in Figure 1, upon generating the initial answer to the given question, an additional feedback instruction is utilized to guide ChatGPT in generating feedback information regarding the initial answer. This information contains an evaluation of the correctness of the initial answer. A subsequent modification instruction is employed to alter or refine the initial answer based on the feedback received.
@@
Nevertheless, small models typically lack self-awareness (Weng et al. 2022) and tend to exhibit greater confidence in their generated responses. Consequently, they struggle to assess the quality of their generated outcomes. The capability for self-verification serves as a prerequisite for achieving self-correction. Furthermore, the manner in which self-correction is achieved through multi-step prompt engineering within LMs differs from the spontaneous and one-time correction observed in humans.
@@
  • To the best of our knowledge, we are the first to demonstrate that small language models with even 6 billion parameters possess the capacity for self-correction during response generation without relying on ground truth.
@@
Instruction fine-tuning LMs often struggle to maintain alignment well with human intent after pre-training. However, this issue can be alleviated by employing techniques such as Instruction Fine-Tuning (IFT) (Stiennon et al. 2020; Ouyang et al. 2022) and Reinforcement Learning from Human or AI Feedback (RLHF/RLAIF) (Gao, Schulman, and Hilton 2023; Bai et al. 2022). IFT involves training an LM to generate responses based on provided prompts, which might optionally include task instructions. The data used for fine-tuning is usually less than the original pre-training corpus. These datasets used in IFT are often curated through crowd-sourcing or extracted from an LM that is already capable of generating instruction-following examples. Studies have shown that IFT can enhance zero-shot generalization to various unseen tasks. To optimize the fine-tuning process, various strategies have been proposed to make it more parameter-efficient. These strategies include using adapters (Hu et al. 2021), prompt tuning (Li and Liang 2021), and other techniques.
Chain-of thought in language model. Chain-of-Thought (COT), as an alternative prompt strategy, adopts a chained reasoning approach, incorporating a multi-step reasoning path before generating the final answer. This strategy aims to enhance the model’s capacity to handle complex and multi-step queries. Kojima et al. (2022) introduced a novel yet simple approach to prompt the LM by using the phase Let’s think step by step, which enables reasoning generation in a zero-shot manner. Zhou et al. (2023) further decomposed the questions into multiple sub-questions, encouraging the model to engage in a sequential thought process while arriving at the answer. Overall, COT aims to enhance the answer quality through a multi-step reasoning process.
@@
Self-improvement in language model. Richer and more detailed feedback plays a crucial role in aligning the output of the LM with the user’s preferences and in significantly improving the overall performance of the model. Welleck et al. (2022) introduced a corrector model used to improve the initial response by learning the specific mistakes made by the LM and the appropriate method for modifying them. Similarly, Paul et al. (2023) introduced a critic model to provide feedback on the LM’s reasoning errors, evaluating the generated intermediate steps. The feedback, together with the original question and preceding intermediate reasoning steps, was subsequently fed back to the LM to improve the generation of the next step.
Scaling model size can increase model’s various capabilities(Kaplan et al. 2020), including self-verification (or self-feedback). Recent research efforts have focused on exploring large LMs that employ self-feedback rather than relying on additional models. Madaan et al. (2023) introduced an iterative self-refinement algorithm that alternates between feedback and refinement. They guided the model to provide feedback about the initial response by employing few-shot prompts. This feedback is then passed back to the same model to help refine the initial response. Some methods used self-improvement to improve the reasoning ability of LM without supervised data. For example, the LM first generates multiple COT reasoning paths and corresponding answers for each question (Wei et al. 2022b). The answer with the highest consistency is selected by majority voting. The COT reasoning paths exhibiting high-confidence, along with the selected answers, are augmented by mixed formats to serve as the training data used for IFT. Additionally, the LM verified its multiple generated answers in turn and calculated the verification score based on the number of predicted masked values (Weng et al. 2022). Instead of requesting the LM to answer the same query multiple times, a more effective approach is to prompt the LM to rephrase the query and then answer each rephrased question individually.
@@
($A_{1}^{1}$) Therefore, the correct answer is D. amusement park
@@ Methods
Additionally, we use two methods to enhance the model’s understanding of various self-correction instructions. First, we use gpt-3.5-turbo to generate diverse prompts for different types of instructions mentioned above. We present a brief prompt template below I want you act as a Prompt Rewriter. Your objective is to rewrite a given prompt to make language model to understand. The rewritten prompt must ensure that the requirement remains unchanged. #Given Prompt#:$\left[TaskP\mid COTP\mid SCP\right]$. Second, we randomly select one instruction from TaskP, COTP and SCP sets and combine them. The order of these instructions can also be rearranged, but TaskP is positioned at the beginning.
@@
Answer preparation. To ensure answer diversity, we utilize nucleus sampling (Holtzman et al. 2020) to generate multiple answers. After generating multiple answers for a question, the next step is to evaluate the accuracy of each answer by comparing it with the provided ground truth for the each question. In the case of multiple-choice questions, we extract the options of the final answer through string matching and then directly compare them with the standard answer to check the accuracy.
@@
In our paper, we apply this loss calculation method to good cases. However, it is not suitable for bad cases due to two reasons. First, these bad cases contain error information, which increases the likelihood of the common issue of hallucination inherent in language models. Second, deliberately training the model to first generate incorrect answers and then correcting them does not align with our intention of teaching the LM how to self-correct. Our aim is instead to enable the model to spontaneously correct itself when it generates information inconsistent with its internal knowledge. Therefore, for bad cases, we refrain from computing the loss corresponding to the incorrect answer part in the output during training. Instead, we calculate the loss from the output on self-verification. It means that for bad cases, only the self-verification and modified correct answer contribute to the loss calculation and parameter update. We call this training method Partial Answer Masking (PAM). In Table 1, the underline part in the output is excluded from the loss calculation.
@@
Base language models. Our goal is to evaluate whether we can improve the performance of any LMs through our proposed ISC, even to small LMs. We opt for open-source models with parameters ranging between 6 billion and 13 billion. We use CuteGPT-7B, CuteGPT-13B [^3] https://github.com/Abbey4799/CuteGPT/, ChatGLM-6B [^4] https://github.com/THUDM/ChatGLM-6B, Llama2-7B [^5] https://huggingface.co/meta-llama/Llama-2-7b-hf, Vicuna-7B [^6] https://huggingface.co/lmsys/vicuna-7b-v1.3 and Vicuna-13B [^7] https://huggingface.co/lmsys/vicuna-13b-v1.3 as our instruction fine-tuning base models for continuing training. These models have been fine-tuned on the instruction data and have the ability to follow instructions. CuteGPT is based on the original Llama model structure, expands the Chinese vocabulary and performs pre-training. CuteGPT has two public versions: CuteGPT-7B and CuteGPT-13B. ChatGLM-6B also is an open bilingual language model. It is trained for about 1T tokens of Chinese and English corpus, and demonstrates outstanding performance among language models of equal parameter size. Llama2 is a family of state-of-the-art open-access large language models released by Meta, ranging in scale from 7 billion to 70 billion parameters. Vicuna is also an open-source chatbot trained by fine-tuning Llama.
@@
Besides, in our experiments, to assess the capability for self-modification of LM, we use R2R, R2W, W2R, and W2W to denote the times that the LM modifies the correct answer to the right answer, changes the right answer to the wrong answer, changes the wrong answer to another incorrect answer, and modifies the wrong answer to the right answer, respectively.
@@
Discovery 4: By examining the values within the R2R column, we discern that even though our experiments primarily focused on self-verification of the final answers rather than the problem-solving process, we note that the models, through self-correction, could identify inadequacies in the analytical process and subsequently provide supplementary analysis.
@@
We discover that ISC remains effective for the new task. After the second round of correction, the accuracy of all LMs improve.
@@ Conclusion
We introduce Intrinsic Self-Correction (ISC) in LMs, an approach that utilizes models’ own capabilities to identify and further modify their initial responses autonomously. This strong capability can even be applied to smaller LMs. We first devise a general process for constructing self-correction data, applicable to generating diverse self-correction task training data. Furthermore, we introduce a novel fine-tuning method named PAM to instruct LMs to self-correct. We conduct experiments on several open-source LMs to validate the efficacy of ISC. The experimental results on two distinct tasks consistently demonstrate that the utilization of ISC empowers the models with the capability for self-correction, and improves the accuracy of generated answers. In the best case, the accuracy enhancement reaches up to 5.6%. We also conduct a comprehensive analysis of the ability of the base model to self-validate and self-correct after using ISC. These findings help us understand how ISC works better.
@@ Acknowledgements
Yanghua Xiao is also a member of Research Group of Computational and AI Communication at Institute for Global Communications and Integrated Media, Fudan University.